Status Report

SOHO Daily Meeting Minutes for Thursday, 21 February 2002

By SpaceRef Editor
February 21, 2002
Filed under , ,

* HESSI target: AR 9830

* CMS testing on Friday at 11am during the D/L only station.

* SPWG Friday at 2pm in the EAF

* The next SSR dump will be Sunday

* Commanding schedule for next week:

Monday – D66 until 18:30 UT then D27.
Tuesday – D66 until 18:30 UT then D27.
Wednesday – D27 from 16-18:30 UT then D16.
Thursday – D66 until 18 UT then D27.
Friday – D66 until 16 UT then downlink only station
until 22:30 UT when we have D27.
Saturday – Commanding is good all day.
Sunday – D66 until 18:30 UT then D27.

Spacecraft Status:
SVM: Nominal, Corrective Actions for Therm Mon, TR Therm Mon, and
EXP LCL disabled
CTOF Recovery in progress

Spacecraft Anomalies: None

Accomplished Activities: VIRGO, CELIAS (3, 1 STOF, 2 CTOF),
New Tracking Star (6.95)

Planned Activities: VIRGO, CELIAS, NOM HGA

Upcoming Operations: None

Ground Anomalies:
051/1903 7.5 hours degraded TM. D27. Reason unknown. Recoverable from CDR.
051/2334 2.6 minutes total TM dropout between 233349 – 233713. D27. Possible
RFI. Nonrecoverable.
052/0250 1.3 minutes TM dropout. D46. TCA channel problem. Nonrecoverable.
052/0832 4 minutes commanding delay. D46. CPA aborting commmands.

02/20 DOY 051 – 02/21 DOY 052

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:

CDS: Nominal. JOP152 on AR 9825. Tomorrow: JOP152 on AR 9830.

UVCS: Nominal. Star observations then synoptics. Tomorrow: coronal
hole observations and synoptic.

LASCO: Nominal. C2/C3 synoptics.

EIT: Nominal. 195 Angstrom CME watch and synoptics, supporting JOP152.

MDI: Nominal. Full disk magnetograms and dopplergrams for continuous

TRACE: Nominal. JOP152 on AR 9830 in 195A.

SpaceRef staff editor.